# Consistency of EFT illuminated via relative entropy: A case study in scalar field theory

### v1

@@ 2.2.1 Scalar field theory comprising one external background field or light field and one massive field
    where $\partial_{I}$ for $I=1,2,3,4$ is a derivative in Euclidean space, and $a_{k}$ is a function comprising $\phi_{1}$ and its derivative. We proceed with redefining $\phi_{2}$ (refer to Appendix A for details) and considering it as follows:
@@
    To explicitly express the order of the perturbation expansion about the interaction (15), we introduce the parameter $\lambda$ and consider the following Lagrangian instead of eq. (13).
@@
    Regarding the classical solution at $\lambda$ higher than the second order, please refer to Appendix B. After substituting eq. (22) into eq. (20) and expanding up to the third order of $\lambda$, the resulting expression is as follows:
@@
    Note that the main perturbative correction to $W_{\lambda}$ is quadratic in $\lambda$ because of eq. (12). Regarding the effective action up to the higher order of $\lambda$ than the third order, look at Appendix B. The effective action $W_{\lambda}$ corresponds to the low-energy EFT of $I_{\lambda}$, establishing a connection between the relative entropy and the EFT through eq. (7). From eq. (20), the third term of eq. (7) becomes,
@@
    where $\widetilde{\phi}_{2,0}=0$ was utilized. Note here that eq. (25) is derived without relying on perturbative calculations.
    Now, we elucidate the assumptions in calculating the effective action as fundamental components of the relative entropy. In the aforementioned perturbative calculations of the EFT $W_{\lambda}$, it was assumed that the interaction $I_{\rm I}$ is small and that the following two conditions are upheld: {itembox}[c] Weakly coupled theory
    These two assumptions imply that the interaction between heavy and light fields is weak, and consequently, we refer to theories that satisfy these assumptions as weakly coupled theories. In contrast, throughout this work, theories that do not satisfy one of the two assumptions are referred to as strongly coupled theories. The first condition above is often assumed in the context of perturbative unitarity of $S$-matrix, and exploring parameter regions of theory satisfying its validity is of significant importance, e.g., in predicting the SM Higgs mass Lee:1977eg ; PhysRevLett.38.883 . In eqs. (23) and (24), the effective action for the perturbation of the interaction $I_{\rm I}$ was calculated by substituting the classical solutions back into the action. It is well-known that perturbations about the interaction arise during the calculation of scattering amplitudes, and the effective action can be obtained by matching these scattering amplitudes. Based on this perspective, the tree-level approximation utilized in perturbative unitarity within the $S$-matrix is similar to the tree-level approximation applied in the EFT above calculations. On the other hand, the second condition represents the validity of operator expansion, meaning the validity of the EFT description in low-energy regions. If the operator expansion is invalid, the phenomenon cannot be adequately described without adding an infinite number of operators. The determination of the Wilson coefficients of an infinite number of operators requires a UV theory that corresponds to the EFT. We thus regard any theory for which the operator expansion is invalid as unsuitable for the EFT description.
@@
    where in the first line, we applied $(\partial W_{\lambda}/\partial\lambda)_{\lambda=0}=0$ following eq. (25). From eq. (26), the second order of $\lambda$ in the relative entropy, i.e., the main order of the perturbation expansion, is generally non-negative in low-energy domains. In contrast, the third order of $\lambda$ in the relative entropy can attain a negative value, raising the possibility of violating the non-negativity property in the perturbative calculation. Even for tertiary and subsequent $\lambda$ contributions, the non-negativity of relative entropy may be disrupted, and the sign of relative entropy may vary based on the perturbation order (see Sec. 3). Since the relative entropy calculated without approximation satisfies the non-negativity (2), the right-hand side of eq. (26) should be non-negative if the perturbation is valid. Based on the facts above, it can be inferred that the sign of the perturbatively calculated relative entropy serves as a potential indicator of the effectiveness of the perturbation of the interaction $\lambda I_{\rm I}$. In short, theories that violate the non-negativity of relative entropy are strongly coupled, rendering the perturbative description of physics invalid. In Sec. 3, two scalar field theories are analyzed as examples, and the validity of perturbations is explored in more detail through the perturbative calculation of the relative entropy, followed by a comparison with the non-perturbative calculation results.
@@
    where $W_{\lambda}[\widetilde{\phi}_{1,\lambda}]=W_{\lambda}[\widetilde{\phi}_{1,0}]+\mathcal{O}(\lambda^{4})$ was used in conjunction with $\widetilde{\phi}_{1,\lambda}=\widetilde{\phi}_{1,0}+\mathcal{O}(\lambda^{2})$. Note that, according to eq. (12), the main perturbative correction to $W_{\lambda}$ is the second order of $\lambda$. The third term of eq. (7) can be obtained from eq. (29) and $\widetilde{\phi}_{2,0}=0$ as,
@@
    where, in the first line, $(\partial W_{\lambda}/\partial\lambda)_{\lambda=0}=0$ was used following eq. (32). Equation (34) is equivalent to eq. (26) up to the third order of $\lambda$ by substituting $\phi_{1}$ with $\widetilde{\phi}_{1,0}$. Thus, it becomes evident that up to the low orders of perturbation, regardless of whether $\phi_{1}$ is the dynamical or background field, the method and results of the calculation remain the same, with the only difference being the replacement of $\phi_{1}\to\widetilde{\phi}_{1,0}$. It should be mentioned that the $\lambda$ dependence of $\widetilde{\phi}_{1,\lambda}$ can be crucial in the higher-order perturbative calculations.
@@ 2.2.2 $N$ scalar fields theory with one heavy field
To make the order of the perturbation expansion explicit, we introduce the parameter $\lambda$ and define the following Lagrangian:
@@
$\left(\frac{\partial W_{\lambda}^{(N)}}{\partial\lambda}\right)_{\lambda=0}=\frac{\partial}{\partial\lambda}I_{\lambda}^{(N)}[\widetilde{\phi}_{1,0}^{(N)},\cdots,\widetilde{\phi}_{N-1,0}^{(N)},\widetilde{\phi}_{N,0}^{(N)}]=0.$ (47)
From eqs. (7), (45), (46), and (47), the relative entropy of the weakly coupled theory, up to the third order of $\lambda$, is obtained as,
@@ 2.2.3 $N$ scalar fields theory with $N-L$ heavy fields
Given the two Secs. 2.2.1 and 2.2.2 above, we will analyze the more general case. In the concrete systems of the following sections, the formulas in this case will not be applied, but it may be worthwhile to consider this case to derive bounds for more general theories. We are examining the identical Lagrangian as outlined in eq. (35). However, we are assuming that $\vec{\Phi}_{N-L}:=(\phi_{L+1},\cdots,\phi_{N})$ represents a set of heavy real scalar fields, while $\vec{\phi}_{L}:=(\phi_{1},\cdots,\phi_{L})$ corresponds to a set of real scalar fields associated with particles having masses smaller than $\vec{\Phi}_{N-L}$’s mass. Under these assumptions, the Lagrangian in eq. (35) can be expressed as follows:
@@
where, from eq. (51), it was used that there is no linear term of $\phi_{n}$ in ~$\widetilde{\mathcal{L}}^{(N)}_{0}$ and that $\left(\partial\widetilde{W}_{\lambda}^{(N)}/\partial\lambda\right)_{\lambda=0}=0$ holds. Also, this inequality is a generalization of eq. (33).
@@ 3 Relative entropy in an effective field theory of inflation
where $M_{\rm Pl}$ represents the reduced Planck mass, $R$ is the Ricci scalar, $\phi$ is an inflaton field, $X:=-\frac{1}{2}(\partial_{\mu}\phi)^{2}$, $c_{j}$ is the Wilson coefficient of the operator $X^{j}$, and $V_{\rm soft}(\phi)$ is assumed to softly break a shift symmetry of $\phi$111 As discussed later, we assume that the soft-breaking term $V_{\rm soft}$ does not affect hard processes associated with heavy fields. We thus refer to $V_{\rm soft}$ as the soft-breaking term. . In general, EFT-based analyses of physical phenomena offer the advantage of making quantitative predictions without requiring associated UV theories that generate the EFT. It is, therefore, essential to identify parameter regions where the EFT loses its ability to predict physical phenomena quantitatively (where the UV theory generating the EFT becomes indispensable) to ensure the reliability of EFT-based predictions of phenomena. Furthermore, the existence of a UV theory corresponding to the EFT is essential for attributing significance to the EFT predictions. We assume the validity of the EFT description in eq. (65), predicated on the validity of the operator $X$ expansion, which refers to the quantitative truncation of higher-dimensional operators at any finite order $n\geq 1$ (also see Sec. 5 for discussions of this assumption). If that assumption is violated, an infinite number of higher-dimensional operators could be associated with the correct physical description. This implies that the UV theory generating eq. (65) is essential for describing physical phenomena.
@@ 3.1 Two scalar fields theory
where $\phi_{1}$ and $\phi_{2}$ refer to real scalar fields, while $V(\phi_{i})$ denotes a potential. To maintain the flatness of the inflaton potential, we postulate an approximate U(1) symmetry under a phase shift of $\phi_{1}+i\phi_{2}:=re^{i\theta}$ for the potential $V(\phi_{i})$. The action (72) is reformulated by incorporating the fields $r$ and $\theta$ in the following manner:
@@
where $V_{r}(r)$ remains invariant under a constant shift of $\theta$, while $V_{\text{soft}}(\theta)$ introduces a soft break in the shift symmetry and does not impact hard processes involving the heavy field $r$. In the following analysis, the EFT of inflation is examined, originating from theories featuring two typical potentials: a U(1) symmetric potential and a U(1) symmetry-breaking potential. The focus of the analysis is to evaluate the conditions under which the EFT description remains valid.
@@ 3.1.1 U(1) symmetric potential
where, in eq. (77), higher derivative terms of $\phi$ vanish for the inflationary background $\phi=\phi_{0}$ due to the slow-roll approximation. Please refer to Appendix B for instructions on calculating eq. (77). In particular, the effective action (77) corresponds to the EFT (65), and the Wilson coefficients in eq. (65) are provided as,
@@
It should be noted here that the relative entropy equals a linear combination of the higher-order operator appearing in the eqs. (77) or (65). From eq. (80), we observe that the theory (75) fully upholds the non-negativity of relative entropy. In contrast, as we will see later, the non-negativity of relative entropy can be violated through the invalidation of perturbation for the U(1) symmetry-breaking potential.
Now, we reconsider the relative entropy for the U(1) symmetric potential up to the full order of $\lambda$ and reconfirm its non-negativity. From eqs. (76) and (21), the classical solution is obtained as,
@@
Therefore, we have observed that the relative entropy remains non-negative for any perturbation order. In this sense, the perturbations of this theory denote a faithful representation of physics, and its U(1) symmetric case would be a weakly coupled theory. In Ref. Kim:2021pbr (also see Sec. 4 of this paper), it was discussed that the perturbative unitarity of this model is satisfied even in high-energy scattering amplitudes. This is qualitatively consistent with the considerations of relative entropy mentioned above, which implicitly encodes the unitarity.
3.1.2 U(1) symmetry-breaking potential
Next, we consider a U(1) symmetry-breaking potential defined as follows:
@@ 3.1.2 U(1) symmetry-breaking potential
where $\tilde{\lambda}_{3}$ and $\tilde{\lambda}_{4}$ are dimensionless parameters characterizing the U(1) symmetry-breaking, and the U(1) symmetry of the potential $V_{r}$ is restored for $\tilde{\lambda}_{3}=\tilde{\lambda}_{4}=1$ (eq. (84) is the same as eq. (74) for $\tilde{\lambda}_{3}=\tilde{\lambda}_{4}=1$). Regarding the parameterization of eq. (84), please also refer to Ref. Kim:2021pbr . From eq. (84), the action (73) is expressed as,
@@
Let us concentrate on the inflationary background and calculate the relative entropy, similar to the U(1) symmetric potential. From eqs. (23), (24), and (86), the Euclidean effective action is calculated, e.g., up to the third order of $\lambda$, as follows:
@@
In Appendix B, we present the effective action up to the sixth order of $\lambda$. Upon combining eqs. (26), (87), and (88), the relative entropy can be derived as,
@@
Here, the relative entropy equals a linear combination of higher dimensional operators appearing in the EFT (87). This is because the difference between theories with and without higher dimensional operators is quantified in terms of the relative entropy. The relative entropy up to the sixth order of $\lambda$ is calculated in Appendix B. The following are the listed results:
@@
Before conducting numerical evaluations of the relative entropy (90), we present an analysis of the behavior of the relative entropy in the large $\tilde{\lambda}_{3}$ region. Let us consider the cases $1\ll|\tilde{\lambda}_{3}|$ and $|\tilde{\lambda}_{4}|\sim 1$. The dominant contribution to $S^{(j)}$ for $j\geq 3$ is then proportional to $(\tilde{\lambda}_{3})^{j-2}$, as illustrated in the Feynman diagrams in Fig. 1. From eqs. (91)-(95), we can further comprehend this behavior. After substituting the background field $(\partial_{\mu}\phi)^{2}=-\dot{\phi}_{0}^{2}<0$, it is important to note that for $0<\tilde{\lambda}_{3}$ and $1\ll|\tilde{\lambda}_{3}|$, the relative entropy is positive at even orders of the perturbations and negative at odd orders of the perturbations. In short, this denotes that the relative entropy lacks a clear sign in the perturbative calculations for the strongly coupled theory.
Figure 1: The dominant Feynman diagrams contributing to $S^{(j)}$ in $1\ll|\tilde{\lambda}_{3}|$ regimes. The diagrams (a), (b), (c), and (d) represent the contributions to $S^{(3)}$, $S^{(4)}$, $S^{(5)}$, and $S^{(6)}$, respectively. The black and the circled cross vertices indicate the interaction proportional to $\tilde{\lambda}_{3}$ and $a_{1}(\phi)$, respectively. For $\tilde{\lambda}_{3}>0$, the odd order of the perturbation in the relative entropy takes negative values while the even ones are positive because of $S^{(j)}\propto(\tilde{\lambda}_{3})^{j-2}(-\dot{\phi}_{0}^{2})^{j}$, where $j>3$. In contrast, for $\tilde{\lambda}_{3}<0$, the non-negativity of relative entropy holds.
We will now proceed to the numerical assessment of the relative entropy. Let us normalize the relative entropy of (90) for $S^{(2)}$ and define the normalized relative entropy up to the $j$-th order of the perturbation as,
@@
where $\lambda=1$ has been selected. In the weakly coupled theory, the perturbations faithfully describe physics, and due to the non-negativity of relative entropy, $s^{(j)}$ should be non-negative regardless of $j$. Due to the non-negativity of relative entropy, eq. (96) should satisfy the following inequalities for the weakly coupled theory, e.g., for $j=3$ and 5,
@@
  • $\tilde{\lambda}_{4}=1$ and $\tilde{\lambda}_{3}=10^{2}$ — We select a positive value for $\tilde{\lambda}_{3}$ to examine the non-negativity of relative entropy and its violation in the perturbative calculations. From eqs. (91)(95), the $j$-th order effect of the perturbation to the normalized relative entropy (96) increases linearly with $(\dot{\phi}_{0}/m_{2}r_{\rm min})^{2j-4}$. In the left panel of Fig. 2, we show the normalized relative entropy up to the $j$-th order as a function of $\dot{\phi}_{0}/m_{2}r_{\rm min}$. The grey line shows $s^{(2)}$, i.e., the main contribution in the perturbative calculation. The red, blue, green, and purple curves represent $s^{(3)}$, $s^{(4)}$, $s^{(5)}$, and $s^{(6)}$, respectively. We also display the numerically evaluated $s^{(\infty)}$ (full order result at the tree-level) as a black dotted curve. In the small $\dot{\phi}_{0}/m_{2}r_{\rm min}$ regions, the colored curves almost coincide with the grey line and black dotted curve, and it is clear that the perturbative calculation is working well. As $\dot{\phi}_{0}/m_{2}r_{\rm min}$ increases, the colored curves deviate from the grey line and the black dotted curve, and the perturbative calculations become progressively invalid. In particular, $s^{(3)}$ (red) almost reaches 10%-100 % of $s^{(2)}$ (grey) and $s^{(\infty)}$(black dotted) around $\dot{\phi}_{0}/m_{2}r_{\rm min}\sim 0.1$, and the non-negativity of relative entropy is broken in the third order of the perturbation. In regions where the value of $\dot{\phi}_{0}/m_{2}r_{\rm min}$ exceeds $\sim 0.1$, the sign of relative entropy changes with each higher order of $\lambda$, rendering the perturbative calculation an inaccurate representation of physics. The variation in the sign of the relative entropy, depending on the order of the perturbation, can also be understood from Fig. 1. The non-negativity of relative entropy is linked to fundamental physics properties, notably unitarity. When the sign of the relative entropy changes, it indicates pathological behavior in the perturbative calculations, and the EFT requires evaluation incorporating all orders of the operator expansion.

### v5

@@ 2.3.1 Scalar field theory comprising one external background field or light field and one massive field
    where $\partial_{I}$ for $I=1,2,3,4$ is a derivative in Euclidean space, and $a_{k}$ is a function comprising $\phi_{1}$ and its derivative. Although it is straightforward to extend the following discussion to cases where higher derivative terms of the heavy fields are present in eq. (11) by including derivative operators in the coefficients $a_{k}(\phi_{1})$, we neglect them since our focus is on low-energy EFTs arising from integrating out heavy fields, where the typical energy scales are smaller than the heavy field masses. We proceed with redefining $\phi_{2}$ (refer to Appendix A for details) and considering it as follows:
@@
    To explicitly express the order of the perturbative expansion about the interaction (15), we introduce the parameter $\lambda$ and consider the following Lagrangian instead of eq. (13).
@@
    Regarding the classical solution at $\lambda$ higher than the second order, please refer to Appendix C. After substituting eq. (22) into eq. (20) and expanding up to the third order of $\lambda$, the resulting expression is as follows:
@@
    Note that the main perturbative correction to $W_{\lambda}$ is quadratic in $\lambda$ because of eq. (12). Regarding the effective action up to the higher order of $\lambda$ than the third order, look at Appendix C. The effective action $W_{\lambda}$ corresponds to the low-energy EFT of $I_{\lambda}$, establishing a connection between the relative entropy and the EFT through eq. (7). From eq. (20), the third term of eq. (7) becomes,
@@
    where $\widetilde{\phi}_{2,0}=0$ was utilized. Note here that eq. (25) is derived under the tree-level approximation without relying on operator expansions.
@@
    where in the first line, we applied $(\partial W_{\lambda}/\partial\lambda)_{\lambda=0}=0$ following eq. (25). From eq. (27), the second order of $\lambda$ in the relative entropy, i.e., the main order of the perturbative expansion, is generally non-negative in low-energy domains. In contrast, the third order of $\lambda$ in the relative entropy can attain a negative value, raising the possibility of violating the non-negativity property in the perturbative calculation. Even for tertiary and subsequent $\lambda$ contributions, the non-negativity of relative entropy may be disrupted, and the sign of relative entropy may vary based on the perturbation order (see Sec. 3). Since the relative entropy calculated without approximation satisfies the non-negativity (2), the right-hand side of eq. (27) should be non-negative if the perturbative calculation is valid. Based on the facts above, it can be inferred that the sign of the perturbatively calculated relative entropy serves as a potential indicator of the effectiveness of the perturbative expansion of the interaction $\lambda I_{\rm I}$. In short, theories that violate the non-negativity of relative entropy are strongly coupled, rendering the perturbative description of physics invalid. In Sec. 3, two scalar field theories are analyzed as examples, and the validity of perturbation theory is explored in more detail through the perturbative calculation of the relative entropy, followed by a comparison with the non-perturbative calculation results.
@@
    where $W_{\lambda}[\widetilde{\phi}_{1,\lambda}]=W_{\lambda}[\widetilde{\phi}_{1,0}]+\mathcal{O}(\lambda^{4})$ was used in conjunction with $\widetilde{\phi}_{1,\lambda}=\widetilde{\phi}_{1,0}+\mathcal{O}(\lambda^{2})$ (see Appendix B). Note that, according to eq. (12) [^5] Here, we consider the Lagrangian (11), which satisfies eq. (12). Then, the classical solution for $\phi_{2}$ is zero at zeroth order in $\lambda$, and the tadpole diagrams associated with $\phi_{2}$ also vanish at zeroth order in $\lambda$. From this, Feynman diagrams contributing to the effective action $W_{\lambda}$ when integrating the heavy field $\phi_{2}$ out cannot appear at the first order of $\lambda$ because the interaction $\lambda I_{I}$ between $\phi_{2}$ and the light fields is the first order of $\lambda$. , the main perturbative correction to $W_{\lambda}$ is the second order of $\lambda$. The third term of eq. (7) can be obtained from eq. (30) and $\widetilde{\phi}_{2,0}=0$ as,
@@
    where, in the first line, $(\partial W_{\lambda}/\partial\lambda)_{\lambda=0}=0$ was used following eq. (33). Equation (35) is equivalent to eq. (27) up to the third order of $\lambda$ by substituting $\phi_{1}$ with $\widetilde{\phi}_{1,0}$. Thus, it becomes evident that up to the low orders of perturbative expansion, regardless of whether $\phi_{1}$ is the dynamical or background field, the method and results of the calculation remain the same, with the only difference being the replacement of $\phi_{1}\to\widetilde{\phi}_{1,0}$. It should be mentioned that the $\lambda$ dependence of $\widetilde{\phi}_{1,\lambda}$ can be crucial in the higher-order perturbative calculations.
@@ 2.3.2 $N$ scalar fields theory with one heavy field
To make the order of the perturbative expansion explicit, we introduce the parameter $\lambda$ and define the following Lagrangian:
@@
$\left(\frac{\partial W_{\lambda}^{(N)}}{\partial\lambda}\right)_{\lambda=0}=0.$ (48)
Note that the partial derivative is performed at $\lambda=0$ (also see Appendix D for the derivation of eq. (48)). From eqs. (7), (46), (47), and (48), the relative entropy of the weakly coupled theory, up to the third order of $\lambda$, is obtained as,
@@ 2.3.3 $N$ scalar fields theory with $N-L$ heavy fields
Given the two Secs. 2.3.1 and 2.3.2 above, we will analyze the more general case. In the concrete systems of the following sections, the formulas in this case will not be applied, but it may be worthwhile to consider this case to derive bounds for more general theories. We are examining the identical Lagrangian as outlined in eq. (36). However, we are assuming that $\vec{\Phi}_{N-L}:=(\phi_{L+1},\cdots,\phi_{N})$ represents a set of heavy real scalar fields, while $\vec{\phi}_{L}:=(\phi_{1},\cdots,\phi_{L})$ corresponds to a set of real scalar fields associated with particles having masses smaller than $\vec{\Phi}_{N-L}$’s mass. Under these assumptions, the Lagrangian in eq. (36) can be expressed as follows:
@@
where, from eq. (52), it was used that there is no linear term of $\phi_{n}$ in ~$\widetilde{\mathcal{L}}^{(N)}_{0}$ and that $\left(\partial\widetilde{W}_{\lambda}^{(N)}/\partial\lambda\right)_{\lambda=0}=0$ holds. Regarding $\left(\partial\widetilde{W}_{\lambda}^{(N)}/\partial\lambda\right)_{\lambda=0}$, it is important to note that the partial derivative is evaluated at $\lambda=0$. Also, this inequality is a generalization of eq. (34).
@@ 3 Relative entropy in an effective field theory of inflation
where $M_{\rm Pl}$ represents the reduced Planck mass, $R$ is the Ricci scalar, $\phi$ is an inflaton field, $X:=-\frac{1}{2}(\partial_{\mu}\phi)^{2}$, $c_{j}$ is the Wilson coefficient of the operator $X^{j}$, and $V_{\rm soft}(\phi)$ is assumed to softly break a shift symmetry of $\phi$666 As discussed later, we assume that the soft-breaking term $V_{\rm soft}$ does not affect hard processes associated with heavy fields. We thus refer to $V_{\rm soft}$ as the soft-breaking term. . In general, EFT-based analyses of physical phenomena offer the advantage of making quantitative predictions without requiring associated UV theories that generate the EFT. It is, therefore, essential to identify parameter regions where the EFT loses its ability to predict physical phenomena quantitatively (where the UV theory generating the EFT becomes indispensable) to ensure the reliability of EFT-based predictions of phenomena. Furthermore, the existence of a UV theory corresponding to the EFT is essential for attributing significance to the EFT predictions. We assume the validity of the EFT description in eq. (66), predicated on the validity of the operator $X$ expansion, which refers to the quantitative truncation of higher-dimensional operators at any finite order $n\geq 1$ (also see Sec. 2.2 for discussions of this assumption). If that assumption is violated, an infinite number of higher-dimensional operators could be associated with the correct physical description. This implies that the UV theory generating eq. (66) is essential for describing physical phenomena.
@@ 3.1 Two scalar fields theory
where $\phi_{1}$ and $\phi_{2}$ refer to real scalar fields, while $V(\phi_{i})$ denotes a potential. To maintain the flatness of the inflaton potential, we postulate an approximate U(1) symmetry under a phase shift of $\phi_{1}+i\phi_{2}:=re^{i\theta}$ for the potential $V(\phi_{i})$. When considering the potential $V(\phi_{i})$ at the renormalizable level, the action (73) is reformulated by incorporating the fields $r$ and $\theta$ in the following manner:
@@
where $V_{r}(r)$ remains invariant under a constant shift of $\theta$, while $V_{\text{soft}}(\theta)$ introduces a soft break in the shift symmetry and does not impact hard processes involving the heavy field $r$. The following analysis examines the EFT of inflation, originating from theories featuring two typical potentials: a U(1) symmetric potential and a modified potential. The focus of the analysis is to evaluate the conditions under which the EFT description remains valid.
@@ 3.1.1 U(1) symmetric potential
where, in eq. (78), higher derivative terms of $\phi$ vanish for the inflationary background $\phi=\phi_{0}$ due to the slow-roll approximation. Please refer to Appendix C for instructions on calculating eq. (78). In particular, the effective action (78) corresponds to the EFT (66), and the Wilson coefficients in eq. (66) are provided as,
@@
It should be noted here that the relative entropy equals a linear combination of the higher-order operator appearing in the eqs. (78) or (66). From eq. (81), we observe that the theory (76) fully upholds the non-negativity of relative entropy. In contrast, as we will see later, the non-negativity of the relative entropy can break down in the modified case of the U(1)-symmetric potential (75), where perturbative calculations become invalid.
Now, we reconsider the relative entropy for the U(1) symmetric potential (75) up to the full order of $\lambda$ and reconfirm its non-negativity. From eqs. (77) and (21), the classical solution is obtained as,
@@
Therefore, we have observed that the relative entropy remains non-negative for any perturbation order. In this sense, the perturbative calculations of this theory denote a faithful representation of physics, and its U(1) symmetric case (75) would be a weakly coupled theory. In Ref. Kim:2021pbr (also see Sec. 4 of this paper), it was discussed that the perturbative unitarity of this model is satisfied even in high-energy scattering amplitudes. This is qualitatively consistent with the considerations of relative entropy mentioned above, which implicitly encodes the unitarity.
3.1.2 A modified version of potential in eq. (75)
Next, we consider a modified version of the potential in eq. (75) at the renormalizable level defined as follows:
@@
where $\tilde{\lambda}_{3}$ and $\tilde{\lambda}_{4}$ are dimensionless parameters characterizing the deviations from eq. (75), and eq. (85) is the same as eq. (75) for $\tilde{\lambda}_{3}=\tilde{\lambda}_{4}=1$. This work focuses on the renormalizable potential in eq. (73), where $\phi_{1}+i\phi_{2}=re^{i\theta}$. Regarding the parameterization of eq. (85), please also refer to Ref. Kim:2021pbr . From eq. (85), the action (74) is expressed as,
@@
Let us concentrate on the inflationary background and calculate the relative entropy, similar to the U(1) symmetric potential in subsubsection 3.1.1. From eqs. (23), (24), and (87), the Euclidean effective action is calculated, e.g., up to the third order of $\lambda$, as follows:
@@
In Appendix C, we present the effective action up to the sixth order of $\lambda$. Upon combining eqs. (27), (88), and (89), the relative entropy can be derived as,
@@
Here, the relative entropy equals a linear combination of higher dimensional operators appearing in the EFT (88). This is because the difference between theories with and without higher dimensional operators is quantified in terms of the relative entropy. The relative entropy up to the sixth order of $\lambda$ is calculated in Appendix C. The following are the listed results:
@@
Before conducting numerical evaluations of the relative entropy (91), we present an analysis of the behavior of the relative entropy in the large $\tilde{\lambda}_{3}$ region. Let us consider the cases $1\ll|\tilde{\lambda}_{3}|$ and $|\tilde{\lambda}_{4}|\sim 1$. The dominant contribution to $S^{(j)}$ for $j\geq 3$ is then proportional to $(\tilde{\lambda}_{3})^{j-2}$, as illustrated in the Feynman diagrams in Fig. 1. From eqs. (92)-(96), we can further comprehend this behavior. After substituting the background field $(\partial_{\mu}\phi)^{2}=-\dot{\phi}_{0}^{2}<0$, it is important to note that for $0<\tilde{\lambda}_{3}$ and $1\ll|\tilde{\lambda}_{3}|$, the relative entropy is positive at even orders of the perturbative expansions and negative at odd orders of the perturbative expansions. In short, this denotes that the relative entropy lacks a clear sign in the perturbative calculations for the strongly coupled theory.
Figure 1: The dominant Feynman diagrams contributing to $S^{(j)}$ in $1\ll|\tilde{\lambda}_{3}|$ regimes. The diagrams (a), (b), (c), and (d) represent the contributions to $S^{(3)}$, $S^{(4)}$, $S^{(5)}$, and $S^{(6)}$, respectively. The black and the circled cross vertices indicate the interaction proportional to $\tilde{\lambda}_{3}$ and $a_{1}(\phi)$, respectively. For $\tilde{\lambda}_{3}>0$, the odd order of the perturbative expansion in the relative entropy takes negative values while the even ones are positive because of $S^{(j)}\propto(\tilde{\lambda}_{3})^{j-2}(-\dot{\phi}_{0}^{2})^{j}$, where $j>3$. In contrast, for $\tilde{\lambda}_{3}<0$, the non-negativity of relative entropy holds.
We will now proceed to the numerical assessment of the relative entropy. Let us normalize the relative entropy of (91) for $S^{(2)}$ and define the normalized relative entropy up to the $j$-th order of the perturbative expansion as,
@@
where $\lambda=1$ has been selected. In the weakly coupled theory, the perturbative calculations faithfully describe physics, and due to the non-negativity of relative entropy, $s^{(j)}$ should be non-negative regardless of $j$. Due to the non-negativity of relative entropy, eq. (97) should satisfy the following inequalities for the weakly coupled theory, e.g., for $j=3$ and 5,
@@
  • $\tilde{\lambda}_{4}=1$ and $\tilde{\lambda}_{3}=10^{2}$ — We select a positive value for $\tilde{\lambda}_{3}$ to examine the non-negativity of relative entropy and its violation in the perturbative calculations. From eqs. (92)(96), the $j$-th order effect of the perturbative expansion to the normalized relative entropy (97) increases linearly with $(\dot{\phi}_{0}/m_{2}r_{\rm min})^{2j-4}$. In the left panel of Fig. 2, we show the normalized relative entropy up to the $j$-th order as a function of $\dot{\phi}_{0}/m_{2}r_{\rm min}$. The grey line shows $s^{(2)}$, i.e., the main contribution in the perturbative calculation. The red, blue, green, and purple curves represent $s^{(3)}$, $s^{(4)}$, $s^{(5)}$, and $s^{(6)}$, respectively. We also display the numerically evaluated $s^{(\infty)}$ (full order result at the tree-level) as a black dotted curve. In the small $\dot{\phi}_{0}/m_{2}r_{\rm min}$ regions, the colored curves almost coincide with the grey line and black dotted curve, and it is clear that the perturbative calculation is working well. As $\dot{\phi}_{0}/m_{2}r_{\rm min}$ increases, the colored curves deviate from the grey line and the black dotted curve, and the perturbative calculations become progressively invalid. In particular, $s^{(3)}$ (red) almost reaches 10%-100 % of $s^{(2)}$ (grey) and $s^{(\infty)}$(black dotted) around $\dot{\phi}_{0}/m_{2}r_{\rm min}\sim 0.1$, and the non-negativity of relative entropy is broken in the third order of the perturbative expansion. In regions where the value of $\dot{\phi}_{0}/m_{2}r_{\rm min}$ exceeds $\sim 0.1$, the sign of relative entropy changes with each higher order of $\lambda$, rendering the perturbative calculation an inaccurate representation of physics. The variation in the sign of the relative entropy, depending on the order of the perturbative expansion, can also be understood from Fig. 1. The non-negativity of relative entropy is linked to fundamental physics properties, notably unitarity. When the sign of the relative entropy changes, it indicates pathological behavior in the perturbative calculations, and the EFT requires evaluation incorporating all orders of the operator expansion.